Find the monthly payment needed to amortize a typical $115,000 mortgage loan amortized over 30 years at an annual interest rate of 5.3% compounded monthly. (Round your answers to the nearest cent.) $

Find the total interest paid on the loan. $
Answer #1

Loan amount = $115000, r = 5.3/1200, n = 30*12 = 360 months

\small PMT = \frac{PV*r}{1-(1+r)^{-n}} = \frac{115000*\frac{5.3}{1200}}{1-(1+\frac{5.3}{1200}^{-360})}\approx \$638.6

So, monthly payment is $638.6

Total interest paid = Total payment - loan amount = 638.6 * 360 - 115000 = 114896

So, total interest paid on the loan is $114896.

Please comment if any doubt. Thank you